OpenStudy (anonymous):

Help Please! i am wondering if i got this correct! I will medal. Factor completely: 3x2 + 2x - 1 (1 point) (3x + 1)(x - 1) (3x + 1)(x + 1) (3x - 1)(x + 1) (3x - 1)(x - 1) (I chose A)

OpenStudy (anonymous):

\[3x^2+2x-1\\ as\ 3\times -1=-3, we\ need\ break\ 3\ intwo\ parts\ such\ way\ so\ that\ when\\ subtracting (minus \ sign\ indicates\it) \ these\ parts(factors \ of\ 3 )we\\ get\ 2(coefficient\ of\ x)\ here,3-1=2\\so\ we \ can\ write\ given\ function\ as,\ 3x^2+(3-1)x-1\\ \implies\ 3x^2+3x-x-1\\ \implies\ 3x(x+1)-1(x+1)\\ \implies(x+1)(3x-1)\\ \therefore\ correct\ answer\ is\ (3x-1)(x+1).\]

OpenStudy (anonymous):

sorry Loser66, i had to choose one